# Adaptive Multi-Order Graph Regularized NMF with Dual Sparsity for Hyperspectral Unmixing

### v1

@@ V-A3 Stopping Criteria
For Algorithm 1, two stopping criteria are considered. The first is based on a predefined error tolerance. Once the relative error is less than $\epsilon=10^{-4}$, the iteration stops. The second is the maximum number of iterations, which is set to 3000. The optimization process halts when either of these criteria is satisfied. For Algorithm 2, the maximum number of iterations is 50. The iteration stops when either the maximum number of iterations is reached or the relative error is smaller than $10^{-6}$.
@@ V-A4 Evaluation Metrics
V-B Results on Synthetic Datasets
@@ V-B1 Dataset Description
The Simu-2 dataset is selected from the USGS digital repository [^2] http://speclab.cr.usgs.gov/spectral.lib06, containing reflectance data across 224 spectral channels (0.4-2.5 $\mu$m). This dataset is generated under the LMM framework with four materials, see Fig. 4. The spatial layout includes discrete geometric regions of pure materials and complex mixtures, as shown in Fig. 4.
@@ V-B2 Numerical Results
To emulate real acquisition conditions, the synthetic hyperspectral cubes are corrupted with additive noise calibrated to specific signal-to-noise ratios (SNRs). The SNR quantification follows
@@
[…]
Figure 7: Visualization of the Samson dataset, where (a) 3-D cube and (b) abundances.
Figure 8: Visualization of the Jasper Ridge dataset, where (a) 3-D cube and (b) abundances.
@@ V-C1 Dataset Description
The Samson dataset [^3] http://www.escience.cn/people/feiyunZHU/Dataset_GT.html is extensively used in unmixing, which contains $952\times 952$ pixels, each with 156 spectral channels, covering the wavelength range from 401 nm to 889 nm, and the spectral resolution is up to 3.13 nm. A subregion of $95\times 95$ pixels starting at position $(252,332)$ is selected for analysis. Specifically, this image includes three targets, i.e., Soil, Tree, and Water, as shown in Fig. 7.
The Jasper Ridge dataset [^4] http://www.escience.cn/system/file?fileId=68574 consists of $512\times 614$ pixels, with each pixel recorded across 224 spectral channels covering a wavelength range from 380 nm to 2500 nm, and a spectral resolution of up to 9.46 nm. A subimage of $100\times 100$ pixels is used for analysis. The endmembers include four categories: Tree, Soil, Water, and Road, as shown in Fig. 8.
TABLE III: SAD comparisons for the Samon dataset, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL  MOGNMF
Soil 0.0447 0.0649 0.0129 0.0286 0.0209 0.0166 0.0237 0.0265 0.0094
Tree 0.0434 0.0532 0.0351 0.0716 0.0443 0.0374 0.0419 0.0386 0.0368
Water 0.1045 0.1208 0.1802 0.1389 0.0979 0.1045 0.0977 0.0715 0.0858
Average 0.0642 0.0796 0.0761 0.124 0.0544 0.0528 0.0544 0.0455 0.0447
TABLE IV: SAD comparisons for the Jasper Ridge dataset, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL  MOGNMF
Tree 0.1369 0.1149 0.1370 0.1441 0.0970 0.0967 0.0810 0.1068 0.0776
Soil 0.1367 0.1826 0.1215 0.1328 0.1572 0.1554 0.1187 0.3781 0.1147
Water 0.1066 0.0617 0.0932 0.1052 0.0455 0.0454 0.0543 0.0407 0.0374
Road 0.0559 0.0303 0.0754 0.0664 0.0315 0.0283 0.0335 0.0427 0.0279
Average 0.1090 0.0974 0.1068 0.1121 0.0828 0.0814 0.0696 0.1421 0.0644
Figure 11: Abundance map comparisons for the Samson dataset.
Figure 12: Abundance map comparisons for the Jasper Ridge dataset.
V-C2 Numerical Results
@@
To demonstrate the performance of the proposed MOGNMF on real-world hyperspectral datasets, the SAD between the estimated endmembers and the reference spectra from the USGS spectral library is computed. The SAD values for MOGNMF and other compared methods on the Samson dataset are summarized in Table III. The results show that MOGNMF achieves better performance than competitors. To illustrate the capability of MOGNMF in extracting endmember signatures, Fig. 9 presents a comparison between the extracted signatures and the reference endmembers. The high degree of consistency between them validates the effectiveness of MOGNMF in the HU task. Besides, Table IV shows the resulting SAD obtained by different methods on the Jasper Ridge dataset. It can be observed that our proposed MOGNMF also achieves a lower average SAD value than others. Fig. 10 presents a comparison between the endmember signatures extracted by MOGNMF and the reference endmember signatures, demonstrating that MOGNMF can accurately extract spectral features and effectively distinguish different endmember components.
Furthermore, Fig. 11 presents the abundance maps estimated by various methods on the Samson dataset, where blue represents lower abundance values, and yellow represents higher abundance values, which closely match the overall distribution trend of the reference map. The proposed MOGNMF demonstrates significant advantages in preserving spatial details and more naturally handling boundary transitions. For the Soil components, the proposed MOGNMF excels in retaining the edge details of the Soil regions, providing smoother transitions, and avoiding excessive sparsity or over-concentration in certain areas. In contrast, VCA-FCLS and MVC-NMF show blurring at the boundaries of Soil regions, with noticeable regional errors in abundance distribution, while IRNMF and SGSNMF exhibit overestimation or underestimation. These observations highlight the advantages of the multi-order graph regularization in the proposed MOGNMF. Additionally, Fig. 12 shows the abundance maps estimated by various methods on the Jasper Ridge dataset. The abundance maps obtained by our proposed MOGNMF are highly consistent with the true distribution, effectively recovering the true abundance distribution of the hyperspectral data.
@@ V-D Ablation Studies
In this subsection, ablation experiments are conducted to verify the effectiveness of the multi-order graph and sparsity in our proposed MOGNMF. We first analyze the impact by removing different regularization terms as
@@
Next, we analyze the unmixing results when only using the first-order and second-order nearest neighbor relationships by comparing Case I and the following two models
@@
V-E Discussion
This section discusses the impact of image sizes, impact of endmember numbers, and convergence on the Simu-1 dataset.
V-E1 Impact of Image Sizes
This experiment aims to evaluate the sensitivity to different image sizes, where the SNR is set to 20 dB, the number of endmembers is fixed to 9, and the image sizes are set to $30\times 30$, $40\times 40$, $50\times 50$, $60\times 60$, $80\times 80$, and $100\times 100$. As shown in Fig. 15, as the number of pixels increases, SGSNMF achieves better unmixing performance than SNMF, which is mainly attributed to its integration of spatial group structure and abundance sparsity. Our proposed MOGNMF, by incorporating the multi-order graph and dual sparsity, demonstrates greater adaptability and unmixing capability when handling HSIs with different pixel counts.
@@ V-E2 Impact of Endmember Numbers
V-E3 Convergence Analysis
This study is conducted with 5 endmembers and the SNR ratio of 20. Fig. 17 shows that as the number of iterations increases, the objective function value decreases rapidly and gradually stabilizes. In addition, the relative error versus the number of iterations is provided, demonstrating the good convergence behavior.
@@ VI CONCLUSION
In future work, we are interested in extending it to tensor cases and leveraging neural networks to learn prior structures.

### v2

@@ V-A3 Stopping Criteria
For Algorithm 1, two stopping criteria are considered. The first is based on a predefined error tolerance. Once the relative error is less than $\epsilon_{1}=10^{-4}$, the iteration stops. The second is the maximum number of iterations, which is set to 3000. The optimization process halts when either of these criteria is satisfied. For Algorithm 2, the maximum number of iterations is 50. The iteration stops when either the maximum number of iterations is reached or the relative error is smaller than $\epsilon_{2}=10^{-6}$.
@@ V-A4 Evaluation Metrics
V-B Results on Simulated Datasets
@@ V-B1 Dataset Description
[…]
@@ V-B2 Numerical Results
To emulate real acquisition conditions, the simulated hyperspectral cubes are corrupted with additive noise calibrated to specific signal-to-noise ratios (SNRs). The SNR quantification follows
@@
Tables I and II provide the average SAD and RMSE results on the Simu-1 dataset contaminated by Gaussian white noise under SNR = 10, 20, 30, and $40\,\text{dB}$, respectively, where the top two values are marked as red and blue. It can be observed that when the SNR increases, the unmixing accuracy of all methods improves. Furthermore, compared with HGNMF-FS and RNMF, the proposed MOGNMF demonstrates superior performance under different SNR conditions. This indicates that the introduction of dual sparsity enhances robustness and also highlights the significant advantages of multi-order graph learning.
TABLE I: SAD values on the Simu-1 dataset under different SNR conditions, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL   A2SAN  MOGNMF
SNR=10 0.0968 0.0356 0.0442 0.0627 0.0358 0.0361 0.0438 0.0982 0.0514 0.0320
SNR=20 0.0821 0.0350 0.0524 0.0610 0.0436 0.0340 0.0331 0.0443 0.0421 0.0313
SNR=30 0.0321 0.0109 0.0467 0.0508 0.0468 0.0119 0.0319 0.0320 0.0259 0.0099
SNR=40 0.0122 0.0090 0.0460 0.0149 0.0393 0.0076 0.0267 0.0335 0.0108 0.0063
Average 0.0558 0.0226 0.0473 0.0473 0.0432 0.0224 0.0338 0.0520 0.0326 0.0202
TABLE II: RMSE values on the Simu-1 dataset under different SNR conditions, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL   A2SAN  MOGNMF
SNR=10 0.0943 0.0966 0.0988 0.0932 0.0944 0.0935 0.1180 0.1517 0.0928 0.0899
SNR=20 0.0924 0.0937 0.1008 0.1151 0.0941 0.0955 0.0925 0.1276 0.0981 0.0917
SNR=30 0.0552 0.0310 0.0894 0.0600 0.0599 0.0533 0.0515 0.0394 0.0475 0.0276
SNR=40 0.0391 0.0319 0.0916 0.0392 0.0881 0.0376 0.0619 0.0462 0.0427 0.0219
Average 0.0703 0.0633 0.0952 0.0769 0.0841 0.0700 0.0810 0.0933 0.0683 0.0578
TABLE III: SAD comparisons on the Samon dataset, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL   A2SAN  MOGNMF
Soil 0.0447 0.0649 0.0129 0.0286 0.0209 0.0166 0.0237 0.0265 0.0686 0.0094
Tree 0.0434 0.0532 0.0351 0.0716 0.0443 0.0374 0.0419 0.0386 0.0422 0.0368
Water 0.1045 0.1208 0.1802 0.1389 0.0979 0.1045 0.0977 0.0715 0.0740 0.0858
Average 0.0642 0.0796 0.0761 0.124 0.0544 0.0528 0.0544 0.0455 0.0616 0.0447
TABLE IV: SAD comparisons on the Jasper Ridge dataset, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL   A2SAN  MOGNMF
Tree 0.1369 0.1149 0.1370 0.1441 0.0970 0.0967 0.0810 0.1068 0.0990 0.0776
Soil 0.1367 0.1826 0.1215 0.1328 0.1572 0.1554 0.1187 0.3781 0.2051 0.1147
Water 0.1066 0.0617 0.0932 0.1052 0.0455 0.0454 0.0543 0.0407 0.0914 0.0374
Road 0.0559 0.0303 0.0754 0.0664 0.0315 0.0283 0.0335 0.0427 0.0935 0.0279
Average 0.1090 0.0974 0.1068 0.1121 0.0828 0.0814 0.0696 0.1421 0.1223 0.0644
TABLE V: SAD comparisons on the Urban dataset, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL   A2SAN  MOGNMF
Asphalt 0.1358 0.1490 0.1362 0.1243 0.1152 0.1395 0.1282 0.1699 0.0969 0.1097
Grass 0.1159 0.1083 0.1292 0.1361 0.1266 0.1125 0.1149 0.1281 0.0951 0.1051
Tree 0.0562 0.0837 0.0880 0.0891 0.0481 0.0814 0.0681 0.1248 0.0477 0.0647
Roof 0.1021 0.1255 0.0857 0.1005 0.0686 0.0691 0.0547 0.1057 0.1290 0.0595
Average 0.1025 0.1166 0.1098 0.1125 0.0896 0.0974 0.0915 0.1321 0.0919 0.0848
TABLE VI: SAD values on the Samon dataset under different SNR conditions, where the top two values are marked as red and blue.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL   A2SAN  MOGNMF
SNR=10 0.1338 0.1278 0.1221 0.1127 0.1361 0.1041 0.1007 0.1305 0.1209 0.0964
SNR=20 0.0648 0.0851 0.0636 0.0907 0.0938 0.0595 0.0601 0.0969 0.0864 0.0599
SNR=30 0.0603 0.0620 0.0563 0.0608 0.0619 0.0492 0.0557 0.0750 0.0597 0.0530
SNR=40 0.0596 0.0629 0.0761 0.0521 0.0557 0.0502 0.0497 0.0501 0.0631 0.0464
Average 0.0721 0.0901 0.0795 0.0891 0.0869 0.0658 0.0666 0.0881 0.0825 0.0639
@@
Figure 8: Endmember signatures extracted by MOGNMF on the Urban dataset, where (a) Asphalt, (b) Grass, (c) Tree, and (d) Roof.
Figure 9: Abundance map comparisons on the Samson dataset.
Figure 10: Abundance map comparisons on the Jasper Ridge dataset.
Figure 11: Abundance map comparisons on the Urban dataset.
@@ V-C1 Dataset Description
The Samson dataset [^10] http://www.escience.cn/people/feiyunZHU/Dataset_GT.html is extensively used in unmixing, which contains $952\times 952$ pixels, each with 156 spectral channels, covering the wavelength range from 401 nm to 889 nm, and the spectral resolution is up to 3.13 nm. A subregion of $95\times 95$ pixels starting at position $(252,332)$ is selected for analysis. Specifically, this image includes three targets, i.e., Soil, Tree, and Water.
The Jasper Ridge dataset [^11] http://www.escience.cn/system/file?fileId=68574 consists of $512\times 614$ pixels, with each pixel recorded across 224 spectral channels covering a wavelength range from 380 nm to 2500 nm, and a spectral resolution of up to 9.46 nm. A subimage of $100\times 100$ pixels is used for analysis. The endmembers include four categories: Tree, Soil, Water, and Road.
The Urban dataset consists [^12] https://rslab.ut.ac.ir/data of $307\times 307$ pixels, with each pixel recorded across 210 spectral bands covering a wavelength range from 400 nm to 2500 nm. After removing bands 1–4, 76, 87, 101–111, 136–153, and 198–210 due to dense water vapor absorption and atmospheric effects, 162 bands are retained for analysis. The dataset contains four endmembers: Asphalt, Grass, Tree, and Roof.
@@
V-C2 Numerical Results
To demonstrate the performance of the proposed MOGNMF on real-world hyperspectral datasets, the SAD between the estimated endmembers and the reference spectra from the USGS spectral library is computed. The SAD values for MOGNMF and other compared methods on the Samson dataset are summarized in Table III. The results show that MOGNMF achieves better performance than competitors. To illustrate the capability of MOGNMF in extracting endmember signatures, Fig. 6 presents a comparison between the extracted signatures and the reference endmembers. The high degree of consistency between them validates the effectiveness of MOGNMF in the HU task. Besides, Table IV shows the resulting SAD obtained by different methods on the Jasper Ridge dataset. It can be observed that our proposed MOGNMF also achieves a lower average SAD value than others. Fig. 7 presents a comparison between the endmember signatures extracted by MOGNMF and the reference endmember signatures, demonstrating that MOGNMF can accurately extract spectral features and effectively distinguish different endmember components. Furthermore, to further validate the unmixing capability of MOGNMF on a larger real dataset, the Urban dataset is used. Table V presents the SAD values of different methods on the Urban dataset. The results demonstrate that MOGNMF also performs well on the Urban dataset, and the extracted endmember signatures by MOGNMF, compared with the reference spectra, are shown in Fig. 8, confirming the effectiveness of the proposed method.
Additionally, Fig. 11 presents the abundance maps estimated by various methods on the Samson dataset, where blue represents lower abundance values, and yellow represents higher abundance values, which closely match the overall distribution trend of the reference map. The proposed MOGNMF demonstrates significant advantages in preserving spatial details and more naturally handling boundary transitions. For the Soil components, the proposed MOGNMF excels in retaining the edge details of the Soil regions, providing smoother transitions, and avoiding excessive sparsity or over-concentration in certain areas. In contrast, VCA-FCLS and MVC-NMF show blurring at the boundaries of Soil regions, with noticeable regional errors in abundance distribution, while the abundance maps produced by IRNMF and SGSNMF show both overestimated and underestimated abundance values. These observations highlight the advantages of the multi-order graph regularization in the proposed MOGNMF. Moreover, Fig. 11 shows the abundance maps estimated by various methods on the Jasper Ridge dataset. The abundance maps obtained by our proposed MOGNMF are highly consistent with the true distribution, effectively recovering the true abundance distribution of the hyperspectral data. Similarly, Fig. 11 presents the abundance maps, indicating that MOGNMF also produces abundance results on the Urban dataset that are highly consistent with the ground truth.
@@ V-D Ablation Studies
In this subsection, ablation experiments are conducted to verify the effectiveness of the multi-order graph and sparsity in our proposed MOGNMF. We first analyze the impact by removing different regularization terms as follows.
@@
Next, we analyze the unmixing results when only using the first-order and second-order nearest neighbor relationships by comparing Case I and the following two models.
@@
Figure 14: Parameter sensitivity analysis with respect to $\lambda$ and $\beta$, where (a)-(b) are results when $\alpha=0.1$, (c)-(d) are results when $\alpha=0.01$.
V-E Discussion
This section discusses the parameter analysis, impact of image sizes, impact of endmember numbers, impact of order $K$, convergence analysis, sensitivity analysis, robustness analysis, and convergence analysis.
V-E1 Parameter Analysis
Four critical parameters, namely $\alpha,\beta,\gamma,\lambda$, are considered in this study, where $\gamma$ is estimated according to (38). Therefore, we first analyze the influence of $\beta$ and $\lambda$ while fixing $\alpha=0.1,0.01$. As shown in Fig. 14, under the condition of SNR = 20 and six endmembers, SAD and RMSE exhibit similar variation trends. Specifically, both metrics reach their minimum values when $\lambda$ is within the range of 0.01 to 0.1 and $\beta$ is within the range of 1 to 2. Hence, our experiments are conducted using parameter values within these ranges.
V-E2 Impact of Image Sizes
This experiment aims to evaluate the sensitivity to different image sizes, where the SNR is set to 20 dB, the number of endmembers is fixed to 9, and the image sizes are set to $25\times 25$, $36\times 36$, $49\times 49$, $64\times 64$, $81\times 81$, and $100\times 100$. As shown in Fig. 15, as the number of pixels increases, SGSNMF achieves better unmixing performance than SNMF, which is mainly attributed to its integration of spatial group structure and abundance sparsity. Our proposed MOGNMF, by incorporating the multi-order graph and dual sparsity, demonstrates greater adaptability and unmixing capability when handling HSIs with different pixel counts.
@@ V-E2 Impact of Image Sizes
Figure 17: Impact of order $K$ on the Simu-1 dataset, where (a) SAD and (b) RMSE.
Figure 18: Comparison with different initializations, where (a) SAD and (b) RMSE.
@@
V-E3 Impact of Order $K$
To analyze the influence of graph order $K$ on unmixing performance, Algorithm 2 is evaluated with $K=1,2,3$. Specifically, $K=1$ corresponds to the use of first-order spectral and spatial graphs, $K=2$ includes both first- and second-order graphs, and so on, forming a multi-order graph structure that captures increasingly complex pixel relationships. As shown in Fig. 17, unmixing results improve as the graph order $K$ increases. This improvement can be attributed to the fact that a first-order graph only captures local neighborhood information. When $K=2$, the graph incorporates neighbors of neighbors, and with $K=3$, the information flow within the graph is further expanded across more distant nodes. The use of higher-order graphs thus allows the model to integrate more comprehensive spatial–spectral information and introduces beneficial cross-region redundancy. This extended connectivity contributes to smoother abundance estimations by encouraging consistency among spectrally similar but spatially distant pixels. However, although increasing the graph order $K$ from 1 to 3 leads to significant performance improvements, further increasing the order, such as to $K=4$, brings only marginal gains. This suggests that the structural patterns of higher-order graphs gradually become similar to those of their lower-order counterparts, offering limited additional information. In other words, higher-order graphs beyond the third order do not contribute substantially new or beneficial structural features, as the third-order graph is already sufficient to capture the underlying data structure effectively.
V-E4 Sensitivity Analysis
To investigate the impact of different initialization strategies on the unmixing performance, random initialization and VCA-FCLS initialization are compared. The experiments are conducted on the synthetic dataset with the SNR of 20 dB and six endmembers. As shown in Fig. 18, the VCA-FCLS initialization achieves better unmixing performance compared to the random initialization. This is mainly because VCA can exploit the geometric structure of hyperspectral data to extract potential endmembers, while FCLS can estimate the abundance matrix under nonnegativity and sum-to-one constraints, ensuring physical interpretability.
V-E5 Robustness Analysis
Table VI illustrates the performance of our proposed MOGNMF and other comparison methods on the Samson dataset under different SNR conditions. Since VCA-FCLS does not consider noise or spatial information and heavily relies on the pure pixel assumption, it is sensitive to noise and its performance degrades under low SNR conditions. In contrast, IRNMF, which constructs a group sparse constrained NMF, is more robust to noise than RNMF and achieves relatively good unmixing results under various SNR levels. HGNMF-FS leverages multi-order graph modeling to represent relationships among HSI pixels, which improves its noise resistance. As mentioned above, MOGNMF constructs more accurate multi-order graphs, thereby demonstrating stronger resistance to interference. Overall, our proposed method consistently achieves the best or second-best unmixing results across all SNR conditions, indicating its high accuracy in unmixing and strong robustness under different noise levels.
V-E6 Convergence Analysis
This study is conducted with 9 endmembers and the SNR ratio of 20. Fig. 19 shows that as the number of iterations increases, the objective function value decreases rapidly and gradually stabilizes. In addition, the relative error versus the number of iterations is provided, demonstrating good convergence behavior.
@@ VI Conclusion
However, introducing high-order graph relationships inevitably increases computational and memory overhead. In the future, we are interested in developing lightweight network architectures to reduce complexity while maintaining accuracy.